OpenStudy (anonymous):

The scientific notation 3.14 x 10 ^-3 has what value?

sam (.sam.):



OpenStudy (unklerhaukus):

\[3.14 \times 10 ^{-3}=\frac{3.14}{10^3}=\frac{3.14}{10\times10\times10}=\frac{0.314}{10\times10}=\frac{0.0314}{10}=0.00314\]